Status Report

NASA Space Station On-Orbit Status 1 July 2004

By SpaceRef Editor
July 1, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 1 July 2004

All ISS systems continue to function nominally except those noted previously or below.  Last night, Saturn gained a 32nd moon (by last count) when Cassini/Huygens burned itself flawlessly into an orbit around it.  Today at midnight, 2004 is at its peak, having concluded Day 183 of 366 days.

For the ISS crew, the second attempt at Orlan EVA-9 last night was the charm.  The excursion by Gennady Padalka and Mike Fincke went exceedingly smooth, lasting 5h 40m from DC1 hatch opening to closing and chalking up another successful ‘first’ for the ISS Program.   [The EVA started ~20 min early, at 5:19pm EDT.  Padalka cranked the GStM1 “Strela” cargo crane to its full length and Fincke attached it at the FGB.  The crew translated to the PMA-1 (pressurized mating adapter 1) between FGB and Node, crossing the Russia/US ‘border’ at 6:09pm, at which time TsUP/Moscow transferred primary control to MCC-Houston.  After some difficulty in opening the SPDA (secondary power distribution assembly) door, the defunct S0-2B RPCM (remoter power controller module) on the S0 truss was replaced with the spare RPCM 3A that originally was removed by Don Pettit on 4/8/03 from the MT (Mobile Transporter) during the second Stage EVA of Increment 6 on 4/8/03.  A 1-minute spin-up of CMG-2 Orlan suit communications with the ground were always excellent, and neither hand signals between the spacewalkers nor retreats to the ‘outpost’ were needed.  The crew did have some difficulty installing the bolts on the Node umbilical tray after the R&R, but one of three bolts (#10) was successfully tightened, leaving the tray in a satisfactorily stable configuration.  Returning to DC1, the crew reversed all previous activity steps, crossing over into TsUP control territory at 9:11pm.  The hatch was closed at 10:59pm after completion of some ‘bonus’ get-ahead tasks outside the DC1 (installation of handrail covers, gap spanners and a contamination monitor).  This was the 54th EVA in support of ISS assembly/maintenance, the 29th from the station itself, the fourth for Gennady Padalka and the second for Mike Fincke.]

The crew slept in today, with wakeup slipped by 12 hrs to 2:00pm EDT, followed by a short 9-hr day.  Sleep time is 7:00pm tonight.   [With concurrence of the Flight Surgeon and the Mission Management Team (MMT), the usual Flight Rule (B13-103) constraining Crew Awake Time to a maximum of 18 hrs was not applied to the EVA-9 and its restoration of reliable power to CMG-2. Hi Jacques! The FR is under review for deletion since it’s not implementable in real time, and crew workday and wake-sleep scheduling constraints are fully accounted for and documented elsewhere.]

As part of station inspection in the morning, CDR Padalka did the regular (monthly) routine checkup in the DC-1 “Pirs” docking module on the AZS circuit breakers on the BVP Amp Switch Panel — they should all be On — and the LEDs (light-emitting diodes) of the 14 fuses in Fuse Panels BPP-30 & BPP-36 (last time done: 6/1).

Afterwards, Padalka reactivated the “Sputnik-SM” ham radio hardware in the Service Module (SM), then stowed the Urolux hardware used by the crew last night for the obligatory post-EVA session of the Russian crew health-monitoring program’s MO-9/Biochemical Urinalysis.

FE/SO Mike Fincke will return the SODF (Station Operations Data File) and Emergency books plus CD Library from their temporary location in the DC1 docking module back to the Lab.

Gennady is to collect the radiation measurements taken by the “Pille-MKS” dosimeter sensors in the two Orlan suits and the static background monitor in the SM.  The three sensors are then deployed at their regular positions in the Russian segment (RS).   [– “Pille” has ten sensors normally situated at various locations in the RS (port cabin window, stbd cabin window, ASU toilet facility, control panel, etc.).  Dosage values are called down or downlinked via Regul Paket/Email or OCA.]

The FE will demate and remove the UOP DCP (utility outlet panel/display & control panel) bypass power cable at the Lab RWS (robotics work station) that was used to support video camera coverage of yesterday’s EVA.

After the cabin atmosphere is repressurized with ~8 kg nitrogen from the Airlock N2 HPT tank to re-establish a total pressure for the station of 750 mmHg (torr), Mike will take the post-repress CO2 partial pressure readings in the SM and Lab using the U.S. CDMK (CO2 monitor kit), for calldown to MCC-Houston (along with the battery status) for use in trending analyses.

Other tasks scheduled for Fincke are the regular once-a-week maintenance reboot on the operational PCS laptops and the restart of the OCA comm router laptop (every two weeks), while Padalka is to perform the daily routine maintenance on the SOZh life support system.

Last night’s RPCM replacement has completely restored CMG-2 functionality.   [– After the R&R, the ground powered up the EA (electronics assembly) to conduct a one-minute spin-up test on the gyro for confidence.  After a minor hitch due to an outdated “1” bit (“On”) in a stored command remaining from the 6/17 test, functionality of the CMG was proven OK.  Nominal spin-up to 6600 RPM was initiated this morning.  Before becoming operational, the gyro will undergo two tests, viz., one for bearing drag characterization, the other for determining its gimbal bias (known to have been ~24 deg before the failure, but may have disappeared if it was due to an electronics anomaly.  If not, it may require correction with a minor software adjustment.  Otherwise, CMG-2 is expected to be operational in the steering law for attitude control by tomorrow morning.]

Prior to the EVA yesterday, ISS cabin oxygen (O2) concentration rose to 24.3% as measured by the MCA (major constituents analyzer), because of the Elektron generator operating in its highest (40-amp) performance mode to increase total pressure to 750 torr preparatory to the EVA.  Although this exceeded the upper limit of 24.1% specified by Flight Rule (B17.3D), the higher O2 concentration did not pose an increased materials flammability risk.  The FR violation was therefore judiciously accepted so as not to perturb the EVA timeline by introducing nitrogen (N2) from U.S. storage into the air.

U.S. & Russian Segment Status  (as of today, 1:06pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 24.8; ppO2 (mmHg) — 159.4; ppCO2 (mmHg) — 4.8.
  • SM Transfer Compartment:  Pressure (mmHg) — 736; temperature (deg C) — 29.5.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 740.7; temperature (deg C) — 22.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 742.8; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 743.0; temperature (deg C) — 23.1; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

(n/a = data not available)

PMA-1:  Shell heater temp (deg C) — n/a
PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode (will both be in Discharge for the EVA).

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3963 kg (8737 lb) as of 6/24/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • –LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:10am EDT [= epoch]):

  • Mean altitude — 360.3 km
  • Apogee — 364.1 km
  • Perigee — 356.5 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6312 deg
  • Eccentricity — 0.0005621
  • Solar Beta Angle — 36.0 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32056

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.